# Review of Machine Learning Approaches for Diagnostics and Prognostics of Industrial Systems Using Industrial Open Source Data

### v1

@@ 4 Assessment & Prognosis
This section delves into the health assessment and prognosis problems within PHM, emphasizing their role in evaluating the current health state of industrial machinery and estimating their remaining useful life. Assessment and prognosis are pivotal to implementing predictive maintenance strategies and preemptively addressing system degradation, thereby reducing the incidence of unplanned downtime, extending equipment lifespan, and optimizing maintenance schedules. Similar to the structure of Section 3, we first summarize the challenges in the competition as outlined in Table 5. Subsequently, we detail the specifics of each competition, including the data-driven approaches applied, and provide a consolidated summary of both the problems with their associated datasets and the proposed solutions in Table 6 and Table 7, respectively.
@@ 4.1 Challenges
The pivotal challenge across all competitions in health assessment and prognosis is summarized in Table 5. The common challenge focuses on identifying degradation trends and estimating RUL. In the meantime, diverse industrial scenarios present each competition with unique challenges; for instance, the 2018 PHM NA and 2020 PHM EU are both confronted with variability due to differing operational conditions—sensor discrepancies due to varied tool use in the former, and twelve distinct operational conditions influenced by solid ratios and particle sizes in the latter. Similarly, in 2021 PHM NA, the diverse flight envelopes introduce considerable variability into the data. A common thread in the 2018 and 2021 PHM NA competitions is the critical need for models to robustly classify multiple fault modes to enable precise RUL prediction. Complicating this need is the data imbalance evident in the 2018 PHM NA, where rare fault instances are overshadowed by prevalent healthy data, potentially reducing the models’ capabilities for rare fault detection. In the 2019 PHM NA, the limited availability of ultrasonic signal data, particularly for the validation set, poses a risk of model overfitting, potentially reducing the reliability of data-driven prognostics. The 2020 PHM EU is faced with the uncertainty of unknown faults, as the test set includes operational conditions not represented in the training and validation data, a discrepancy that could lead to diminished model performance. These multifaceted challenges necessitate the development of sophisticated, adaptable machine learning models capable of overcoming data discrepancy, data imbalance, and operational condition variability to deliver accurate and reliable health assessments and prognostic predictions.
@@
Table 5: Overview of Assessment and Prognosis Challenges in PHM Data Competitions
@@
Table 6: Overview of Assessment and Prognosis Problems in PHM Data Competitions
@@
Investigation into the proposed solutions reveals a pronounced inclination towards tree-based methodologies for fault detection and diagnosis. Wu et al. [77], Huang et al. [22], Zhao et al. [79], and the team led by He [19] have each proposed strategies based on random forest algorithms for early degradation mode detection and diagnosis. Extending the exploration of machine learning techniques, Singh et al. [61] evaluated an assortment of models, including Logistic Regression, Generalized Linear Models, Multivariate Adaptive Regression Splines, Support Vector Regression, Random Forest, Multi-Layer Perceptron, and Gradient Boosting. Among these, the Random Forest model distinguished itself with superior performance. Zheng et al. [81], meanwhile, utilized a knowledge distillation approach towards fault detection across various modes, aiming to bolster the detection of infrequent yet critical faults.
Recent advancements in RUL prediction [30] have predominantly hinged on the application of deep-learning algorithms to analyze complex multivariate time series data. Wu et al. [77], Huang et al. [22], and Anqi He and Xiaoning Jin [19] have all utilized Long Short-Term Memory (LSTM) networks and their variations, such as Gated Recurrent Units (GRU) and LSTM-based Metric Regression (LSTM-MR), to capture important features from the raw data. Hsu et al. [20] expanded upon this approach by integrating a Temporal Convolutional Network (TCN) with LSTM, enhanced by an attention mechanism, which facilitated refined feature extraction from sensor data for accurate RUL prediction. Liu et al.’s two-stage deep transfer learning framework aimed at achieving accurate RUL prediction. In the first stage, the developed model leveraged TCN for initial temporal feature learning, followed by domain adversarial learning for data alignment based on one fault mode. Then in the second stage, the first-stage model was fine-tuned based on other fault mode data to handle multiple fault modes and enhance the RUL prediction performance [44]. Distinct from these methods, Zhao et al. [79] explored transformer networks, focusing solely on data from abnormal operation phases, while Lorenti et al. [46] provided a comprehensive comparison of state-of-the-art methods—including TCN, LSTM, attention-based mechanisms, CNN, and Transformer—culminating in the development of the CeRULEo library to facilitate research in the field. Moreover, Singh et al. [61] introduced a novel approach using Dynamic Time Warping (DTW) for RUL estimation, leveraging a library of truncated degradation curves and health score models to refine the final RUL predictions. Such diversity in methodologies not only underscores the complexity of fault diagnosis and prognosis in high-tech manufacturing but also showcases the robustness of machine learning in tackling these intricate challenges.
@@
The 2019 PHM Conference Data Challenge put a spotlight on the intricate task of fatigue crack length estimation and prediction within aluminum structures at different points. The challenge harnessed wave signal data collected by piezoelectric sensors subjected to both static and dynamic tensile stresses. Given that the test dataset was limited to wave signals from several initial loading cycles, participants were challenged to estimate crack lengths where signal data were present and predict future crack growth for specified cycles lacking signal data. The complexity of the problem delineated a clear path: data-driven methodologies were viable when signals existed, whereas scenarios lacking wave data required exploration into physics-based strategies. Reflecting this path, solutions derived from this challenge have predominantly employed a hybrid approach of data-driven and physics-based techniques.
@@
Regarding the estimation of fatigue crack length with accessible wave signals, a variety of data-driven models and feature engineering strategies emerged, both within and beyond the competition’s scope. Karimian and colleagues [28] proposed a neural network architecture reliant on features manually derived from raw signals—such as the Pearson correlation coefficient, phase shifts, energy, and information entropy—to train models for accurate crack length estimation. On a similar trajectory, Kong and associates [32] initiated their approach with signal preprocessing, utilizing techniques like band-pass filtering and phase alignment to mitigate noise and uncertainty before applying physically insightful feature extraction methods. Thereafter, they harnessed a random forest algorithm, optimizing it through feature selection and grid search for hyperparameter fine-tuning, to estimate crack lengths. Consistent with these methods, Youn et al. [78] also leveraged a band-pass filter for feature extraction from raw wave signals, subsequently constructing a Support Vector Regression (SVR) model with hyperparameters optimized via grid search. Rao and collaborators [56], meanwhile, designed an ensemble learning regression model to improve estimation performance with four useful extracted features (root mean square value, correlation coefficient, first peak value, and logarithm of kurtosis).
Predictive modeling for scenarios lacking wave signal data required a pivot towards physics-based techniques, often in conjunction with data-driven insights, to forecast crack progression. Karimian et al. [28] devised a Particle Filter (PF) strategy, integrating the Paris Law and outputs from the previously developed neural network as observational inputs to refine and update the crack propagation pathway. Kong’s team [32] suggested an ensemble prognostics framework under consistent loading conditions, constructing a probability density function (PDF) for each instance. This was followed by a computation of weights derived from each PDF to output the final crack length prediction. Furthermore, when different loading conditions prevailed, Walker’s equation was invoked to forecast crack lengths. Innovatively, Youn et al. [78] introduced a trans-fitting approach, aimed at extracting the crack growth trend from training data and extrapolating it to the test data predictions. Additionally, Rao’s group advanced a variation version of Paris’ Law, aiming to elucidate the correlation between crack progression and the number of loading cycles [56].
4.4 2020 PHM EU (Filtration System)
In the domain of industrial maintenance, filtration systems are pivotal for helping process pollutants from industrial equipment, ensuring seamless system operation. A predominant challenge encountered within these systems is filter clogging - a phenomenon where accumulated pollutants impede flow rates, thereby disrupting standard industrial workflows. Addressing this issue, the PHM-Europe Conference 2020 data challenge concentrated on the prediction of filtration systems’ remaining useful life, crucial for timely maintenance and replacement scheduling. RUL, in this context, is delineated as the time until the pressure differential across the filter breaches a threshold of 20 psi. The challenge provided the PHME20 dataset, collected from a controlled experimental setup designed to simulate filter clogging at various contamination levels. Twelve distinct conditions were established influenced by two operational parameters: solid ratio (%) and particle size (µm). The competition predominantly witnessed the application of traditional machine-learning techniques. The champion solution [45] employed a novel hybrid approach, amalgamating kernel regression with fundamental statistical methodologies. This approach, underscored by complex data analysis, proved efficacious in enhancing the RUL prediction. Meanwhile, the runner-up, Beirami et al. [6], adopted different cutting-edge feature engineering, and machine-learning methods. Their process began with the extraction of features through a rolling window technique and proceeded with feature selection informed by the linear kernel Support Vector Machine (SVM) coefficients, Recursive Feature Elimination (RFE), correlation matrix, and monotonicity testing. A four-layer sequential neural network was their model of choice, supported by K-fold cross-validation throughout the training and validation stages. Capturing the third position, Ince et al. [26] conducted a comparative study of tree-based algorithms and the Bayesian approach. Specifically, random forest and gradient boosting and Gaussian process (GP) regression were utilized to estimate the RUL of the filtration system, supplemented by a novel fault-based RUL assignment which integrated Piecewise RUL_Assignment and Linear RUL_Assignment,
Contrasting with the feature engineering and conventional machine learning strategies, some researchers explored some deep learning methods. Vu et al. [75] introduced a Convolutional Neural Network (CNN)-based deep learning methodology for RUL prediction, marked by two architectural innovations: a Parameterized Fully Connected Layer that adjusts network weights in response to operational parameter shifts, and a multi-head predictor tailored to distinct degradation process stages. Simultaneously, Tian et al. [72] presented a new transfer ensemble learning (TEL) framework, leveraging metric learning alongside domain dissimilarity metric and Kullback–Leibler (KL) divergence, to enhance model generalization from source to target domains. This TEL framework amalgamated with a bidirectional long short-term memory (Bi-LSTM) algorithm, coined as TEL-Bi-LSTM, was offered for RUL estimation under different operating conditions. In another innovative approach, Ince and Genc [25] proposed a joint autoencoder-regression network, a deep neural architecture that fused a convolutional neural network autoencoder with a long-short-term memory network regressor in an end-to-end training paradigm. Genetic algorithms were instrumental in optimizing hyperparameters for this architecture. Additionally, Lee’s team [39] developed a distinct strategy by first establishing a health assessment criterion. They defined a Health Index (HI) for the filter system and utilized K-means clustering for the categorization of the system’s health stages. Subsequent HI predictions were facilitated by the Bi-LSTM algorithm, thus determining the system’s RUL. Lastly, Falcon et al. introduced an innovative sequence modeling technique termed the Neural Turing Machine (NTM) [17]. Conceptualized as a computational architecture, the NTM leverages available data to interact with an external memory component, an approach that facilitates enhanced accuracy in predictions. This model stands out for its ability to generate more precise outcomes when benchmarked against the prevalent LSTM-based solutions that dominate the field.
4.5 2021 PHM NA (Turbofan Engine)
The 2021 North America PHM Data Challenge was primarily focused on the prediction of RUL for turbofan engines [8], specifically under four distinct flight conditions and seven unique failure modes. Participants were required to create predictive models leveraging the N-CMAPSS dataset, aiming to accurately forecast RUL using complex condition monitoring data. The dataset, a collection consisting of 90 synthetic run-to-failure trajectories for training and an additional 38 truncated datasets for testing, served as a comprehensive foundation for developing robust predictive algorithms to predict RUL accurately.
During the competition, the winning team, led by Lovberg, put forward an innovative approach leveraging a deep convolutional neural network (DCNN) [47]. This network was distinguished by its use of dilated convolutions complemented by gated linear unit activations and integration of residual skip connections. Those techniques were designed to expand the network’s receptive field and enhance flexibility, so as to reduce the complexity of the neural network architecture by using less number of parameters, but still having comparable performance. Additionally, they adopted a strategic sequence sampling method, minimizing less informative samples while retaining enough degradation signals for the network’s input. This methodology distinguished their solution and got 1st place in the PHM data challenge competition. Moreover, Solis-Martin et al. and DeVol’s team pursued advancements in DCNNs as well. The former developed a two-level DCNN system where the first-level DCNN focused on extracting vital features from raw data and the second-level DCNN leverages the output from the first level to accurately estimate the remaining useful life [65]. The latter drew upon established deep learning architectures, deploying the basic blocks or modules from the VGG, GoogLeNet, and ResNet designs into their DCNN framework. This enabled a comparative analysis of model performances using a variety of well-known architectural features [14, 15].
Outside of the competition, Cohen et al. [10] addressed concerns regarding the uncertain and poor interpretability of deep learning models by integrating Principal Component Analysis (PCA) to refine time-domain feature sets and subsequently applying four supervised learning techniques, including artificial neural network, random forest, extreme random forest and XGBoost, to estimate RUL. Their innovative use of a custom loss function in conjunction with traditional ANN models got the best results in both Area Under the Receiver Operating Characteristic (AUROC) and Area Under the Precision-Recall (AUPR) metrics. Lastly, the domain of prognostics and health management has witnessed an upsurge in the application of XAI techniques to bolster the interpretability and trustworthiness of machine learning models. Various methods, including Local Interpretable LIME, SHAP, Layer-wise Relevance Propagation (LRP), Image-Specific Class Saliency Maps, and Gradient-weighted Class Activation Mapping (Grad-CAM), have been reported in recent literature [9, 11, 66]. These techniques are increasingly critical in elucidating the decision-making processes of complex models, thereby fostering greater confidence in their predictive capabilities.
@@
Table 7: Overview of Assessment and Prognosis Methodologies in PHM Data Competitions
5 Common Challenges in PHM Data Challenge Competitions
In this section, we will discuss common challenges faced in the PHM data challenge competitions regarding data-related issues (missing data, data imbalance, and domain shift) and model-related issues (model selection, machine learning model interpretability, model robustness and generalization), and analyze relevant solutions applied to solve these challenges.

### v2

@@ 4 Assessment & Prognosis
This section delves into the health assessment and prognosis problems within PHM, emphasizing their role in evaluating the current health state of industrial machinery and estimating their RUL. Assessment and prognosis are pivotal to implementing predictive maintenance strategies and preemptively addressing system degradation, thereby reducing the incidence of unplanned downtime, extending equipment lifespan, and optimizing maintenance schedules. Similar to the structure of Section 3, we first summarize the challenges in the competition as outlined in Table 5. Subsequently, we detail the specifics of each competition, including the data-driven approaches applied, and provide a consolidated summary of both the problems with their associated datasets and the proposed solutions in Table 6 and Table 7, respectively.
@@ 4.1 Challenges
The pivotal challenge across all competitions in health assessment and prognosis is summarized in Table 5. The common challenge focuses on identifying degradation trends and estimating RUL. In the meantime, diverse industrial scenarios present each competition with unique challenges; for instance, the 2018 PHM NA and 2020 PHM EU are both confronted with variability due to differing operational conditions—sensor discrepancies due to varied tool use in the former, and twelve distinct operational conditions influenced by solid ratios and particle sizes in the latter. Similarly, in 2021 PHM NA, the diverse flight envelopes introduce considerable variability into the data. A common thread in the 2018 and 2021 PHM NA competitions is the critical need for models to robustly classify multiple fault modes to enable precise RUL prediction. Complicating this need is the data imbalance evident in the 2018 PHM NA, where rare fault instances are overshadowed by prevalent healthy data, potentially reducing the models’ capabilities for rare fault detection. In the 2019 PHM NA, the limited availability of ultrasonic signal data, particularly for the validation set, poses a risk of model overfitting, potentially reducing the reliability of data-driven prognostics. The 2020 PHM EU is faced with the uncertainty of unknown faults, as the test set includes operational conditions not represented in the training and validation data, a discrepancy that could lead to diminished model performance. These multifaceted challenges necessitate the development of sophisticated, adaptable ML models capable of overcoming data discrepancy, data imbalance, and operational condition variability to deliver accurate and reliable health assessments and prognostic predictions.
Table 5: Overview of Assessment and Prognosis Challenges in PHM Data Challenge Competitions
@@
Table 6: Overview of Assessment and Prognosis Problems in PHM Data Challenge Competitions
@@
Investigation into the proposed solutions reveals a pronounced inclination towards tree-based methodologies for fault detection and diagnosis. [118, 38, 124] and the team led by [34] have each proposed strategies based on random forest algorithms for early degradation mode detection and diagnosis. Extending the exploration of ML techniques, [93] evaluated an assortment of models, including Logistic Regression, Generalized Linear Models, Multivariate Adaptive Regression Splines, Support Vector Regression (SVR), MLP, random forest, etc. Among these, the random forest model distinguished itself with superior performance. Meanwhile, [126] utilized a knowledge distillation approach towards fault detection across various modes, aiming to bolster the detection of infrequent yet critical faults.
Recent advancements in RUL prediction [48] have predominantly hinged on the application of DL algorithms to analyze complex multivariate time series data. [118, 38, 34] have all utilized Long Short-Term Memory (LSTM) networks and their variations, such as Gated Recurrent Units (GRU) and LSTM-based Metric Regression (LSTM-MR), to capture important features from the raw data. [35] expanded upon this approach by integrating a Temporal Convolutional Network (TCN) with LSTM, enhanced by an attention mechanism, which facilitated refined feature extraction from sensor data for accurate RUL prediction. What’s more, Liu et al.’s two-stage deep transfer learning framework aimed at achieving accurate RUL prediction. In the first stage, the developed model leveraged TCN for initial temporal feature learning, followed by domain adversarial learning for data alignment based on one fault mode. Then in the second stage, the first-stage model was fine-tuned based on other fault mode data to handle multiple fault modes and enhance the RUL prediction performance [66]. Distinct from these methods, [124] explored transformer networks, focusing solely on data from abnormal operation phases, while [68] provided a comprehensive comparison of state-of-the-art methods, including TCN, LSTM, attention-based mechanisms, CNN, and Transformer, culminating in the development of the CeRULEo library to facilitate research in the field. Moreover, [93] introduced a novel approach using DTW for RUL estimation, leveraging a library of truncated degradation curves and health score models to refine the final RUL predictions. Such diversity in methodologies not only underscores the complexity of fault diagnosis and prognosis in high-tech manufacturing but also showcases the robustness of ML in tackling these challenges.
@@
The 2019 PHM Conference Data Challenge put a spotlight on the task of fatigue crack length estimation and prediction within aluminum structures at different points. The challenge harnessed wave signal data collected by piezoelectric sensors subjected to both static and dynamic tensile stresses. Given that the test dataset was limited to wave signals from several initial loading cycles, participants were challenged to estimate crack lengths where signal data were present and predict future crack growth for specified cycles lacking signal data. The complexity of the problem delineated a clear path: data-driven methodologies were viable when signals existed, whereas scenarios lacking wave data required exploration into physics-based strategies. Reflecting this path, solutions derived from this challenge have predominantly employed a hybrid approach of data-driven and physics-based techniques.
@@
Regarding the estimation of fatigue crack length with accessible wave signals, a variety of data-driven models and feature engineering strategies emerged, both within and beyond the competition’s scope. [46] proposed a neural network architecture reliant on features manually derived from raw signals, such as the Pearson correlation coefficient, phase shifts, energy, and information entropy, to train models for accurate crack length estimation. Similarly, [50] initiated their approach with signal preprocessing, utilizing techniques like band-pass filtering and phase alignment to mitigate noise and uncertainty before applying physically insightful feature extraction methods. Thereafter, they harnessed a random forest algorithm, optimizing it through feature selection and grid search for hyperparameter fine-tuning, to estimate crack lengths. Consistent with these methods, [121] also leveraged a band-pass filter for feature extraction from raw wave signals, subsequently constructing an SVR model with hyperparameters optimized via grid search. Rao and collaborators, meanwhile, designed an ensemble learning regression model to improve estimation performance with four useful extracted features (root mean square value, correlation coefficient, first peak value, and the logarithm of kurtosis) [86].
Table 7: Overview of Assessment and Prognosis Methodologies in PHM Data Challenge Competitions
@@
Predictive modeling for scenarios lacking wave signal data required a pivot towards physics-based techniques, often in conjunction with data-driven insights, to forecast crack progression. [46] devised a Particle Filter (PF) strategy, integrating the Paris Law and outputs from the previously developed neural network as observational inputs to refine and update the crack propagation pathway. [50] suggested an ensemble prognostics framework under consistent loading conditions, constructing a probability density function (PDF) for each instance. This was followed by a computation of weights derived from each PDF to output the final crack length prediction. Furthermore, when different loading conditions prevailed, Walker’s equation was utilized to forecast crack lengths. Innovatively, [121] introduced a trans-fitting approach, aimed at extracting the crack growth trend from training data and extrapolating it to the test data predictions. Additionally, Rao’s group advanced a variation version of Paris’ Law, aiming to elucidate the correlation between crack progression and the number of loading cycles [86].
4.4 2020 PHM EU (Filtration System)
In the domain of industrial maintenance, filtration systems are pivotal for helping process pollutants from industrial equipment, ensuring seamless system operation. A predominant challenge encountered within these systems is filter clogging - a phenomenon where accumulated pollutants impede flow rates, thereby disrupting standard industrial workflows. Addressing this issue, the PHM-Europe Conference 2020 data challenge concentrated on the prediction of filtration systems’ RUL, crucial for timely maintenance and replacement scheduling. RUL, in this context, is delineated as the time until the pressure differential across the filter breaches a threshold of 20 psi. The challenge provided the PHME20 dataset, collected from a controlled experimental setup designed to simulate filter clogging at various contamination levels. Twelve distinct conditions were established influenced by two operational parameters: solid ratio (%) and particle size (µm).
The champion solution [67] employed a novel hybrid approach, combining kernel regression with fundamental statistical methodologies. This approach, underscored by complex data analysis, proved efficacious in enhancing the RUL prediction. Meanwhile, the runner-up, [10], adopted different cutting-edge feature engineering, and ML methods. Their process began with the extraction of features through a rolling window technique and proceeded with feature selection informed by the linear kernel Support Vector Machine (SVM) coefficients, RFE, correlation matrix, and monotonicity testing. A four-layer sequential neural network was their model of choice, supported by K-fold cross-validation throughout the training and validation stages. Capturing the third position, [43] conducted a comparative study of tree-based algorithms and the Bayesian approach. Specifically, random forest, gradient boosting, and Gaussian process regression were utilized to estimate the RUL of the filtration system, supplemented by a novel fault-based RUL assignment that integrated ”Piecewise RUL Assignment” and ”Linear RUL Assignment”.
Contrasting with the feature engineering and conventional ML strategies, some researchers explored some DL methods. [113] introduced a CNN-based DL methodology for RUL prediction, marked by two architectural innovations: a Parameterized Fully Connected Layer that adjusts network weights in response to operational parameter shifts, and a multi-head predictor tailored to distinct degradation process stages. Additionally, [107] presented a new transfer ensemble learning (TEL) framework, leveraging metric learning alongside domain dissimilarity metric and Kullback–Leibler (KL) divergence, to enhance model generalization from source to target domains. This TEL framework amalgamated with a bidirectional long short-term memory (Bi-LSTM) algorithm, coined as TEL-Bi-LSTM, was offered for RUL estimation under different operating conditions. In another innovative approach, [42] proposed a joint autoencoder-regression network, a deep neural architecture that fused a CNN autoencoder with an LSTM network regressor in an end-to-end training paradigm. Genetic algorithms were instrumental in optimizing hyperparameters for this architecture. Additionally, Lee’s team developed a distinct strategy by first establishing a health assessment criterion [60]. They defined a Health Index (HI) for the filter system and utilized K-means clustering for the categorization of the system’s health stages. Subsequent HI predictions were facilitated by the Bi-LSTM algorithm, thus determining the system’s RUL. Lastly, Falcon et al. introduced an innovative sequence modeling technique termed the Neural Turing Machine (NTM) [30]. Conceptualized as a computational architecture, the NTM leverages available data to interact with an external memory component, an approach that facilitates enhanced accuracy in predictions. This model stands out for its ability to generate more precise outcomes when benchmarked against the prevalent LSTM-based solutions that dominate the field.
4.5 2021 PHM NA (Turbofan Engine)
The 2021 North America PHM Data Challenge was primarily focused on the prediction of RUL for turbofan engines [13], specifically under four distinct flight conditions and seven failure modes. Participants were required to create predictive models leveraging the N-CMAPSS dataset, aiming to accurately forecast RUL using complex condition monitoring data. The dataset, a collection consisting of 90 synthetic run-to-failure trajectories for training and an additional 38 truncated datasets for testing, served as a comprehensive foundation for developing robust predictive algorithms to predict RUL accurately.
During the competition, the winning team, led by Lovberg, put forward an innovative approach leveraging a deep convolutional neural network (DCNN) [69]. This network was distinguished by its use of dilated convolutions complemented by gated linear unit activations and integration of residual skip connections. Those techniques were designed to expand the network’s receptive field and enhance flexibility, so as to reduce the complexity of the neural network architecture by using less number of parameters, but still having comparable performance. Additionally, they adopted a strategic sequence sampling method, minimizing less informative samples while retaining enough degradation signals for the network’s input. This methodology distinguished their solution and got 1st place in the PHM data challenge competition. Moreover, Solis-Martin et al. and DeVol’s team pursued advancements in DCNNs as well. The former developed a two-level DCNN system where the first-level DCNN focused on extracting vital features from raw data and the second-level DCNN leverages the output from the first level to accurately estimate the RUL [97]. The latter drew upon established DL architectures, deploying the basic blocks or modules from the VGG, GoogLeNet, and ResNet designs into their DCNN framework. This enabled a comparative analysis of model performances using a variety of well-known architectural features [24, 25].
Outside of the competition, [19] addressed concerns regarding the uncertain and poor interpretability of deep learning models by integrating principal component analysis (PCA) to refine time-domain feature sets and subsequently applying four supervised learning techniques, including artificial neural network, random forest, extreme random forest and XGBoost, to estimate RUL. Their innovative use of a custom loss function in conjunction with traditional ANN models got the best results in both Area Under the Receiver Operating Characteristic (AUROC) and Area Under the Precision-Recall (AUPR) metrics. Moreover, the domain of PHM has witnessed an upsurge in the application of XAI techniques to bolster the interpretability and trustworthiness of ML models. Various methods, including LIME, SHAP, LRP, Image-Specific Class Saliency Maps, and Gradient weighted Class Activation Mapping (Grad-CAM), have been reported in recent literature [18, 20, 98]. These techniques are increasingly critical in elucidating the decision-making processes of complex models, thereby fostering greater confidence in their predictive capabilities.
5 Challenges and Possible Solutions
By reviewing nine open PHM data challenge competitions over the past six years, these competitions have showcased a variety of ML approaches applied to different industrial systems. In this section, we summarize common challenges regarding data-related issues and model-related issues and analyze relevant solutions applied to solve these challenges. Moreover, we evaluate the limitations of these competitions and suggest future directions for their evolution, aiming to accelerate the development of next-generation ML-driven PHM solutions.
Data-related challenges encompass issues such as missing data, data imbalance, and domain shift while model-related challenges include the critical aspects of model selection, interpretability of ML models, and their robustness and generalization capabilities. These challenges highlight the complexities of developing effective ML methods to solve various PHM problems. The limitations of current PHM competitions reveal a gap in adopting systematic approaches for building effective PHM systems, and a need for multi-modal machine learning analysis.